# Anisotropic anomalous diffusion in microgravity dusty plasma. Part One: Nonequilibrium Statistical Analysis

### v2

@@ 6.2 Histogram Shape in the Cross-field direction
Figure 19: Components of the Bi-q-Gaussian illustrating two distinct populations: a) red $q_{1}=1.3$ and variance $\sigma_{1}=330$, while b) blue $q_{2}=1.63$ and $\sigma_{2}=63$. These are the same as the $q_{v}$ values found for the cross-field velocity distribution in the 70 Pa 0.7 mA case.
The superposition of the two q-Gaussian distributions shown in Fig. 19 d) shows that one diffusion process (in red) is characterized by a small peak but large variance, while the other one (in blue) has both a large peak and large tails. The variance (and the corresponding temperature) of the blue distribution, however, is much smaller. Thus, the observed diffusion in the cross-field direction can be viewed as a superposition of two processes - one with higher temperature close to equilibrium and a second one with a lower temperature but farther from equilibrium.
@@
In Space Physics, it is common to describe the non-Maxwellian velocity distributions using the kappa-distribution which is related to the q-Gaussian by $\kappa=\frac{1}{q-1}$. In [73], three physical mechanisms for generating a Kappa distribution were proposed: (i) Debye shielding, (ii) magnetic field biding, and (iii) temperature fluctuations. It was proposed that $\kappa$ has a negative correlation with the Debye number $N_{D}$ (number of particles in the Debye sphere). As $\kappa$ is inversely related to $q$, we expect a positive correlation between $N_{D}$ and $q$. In other words, inclreasing number of ions in the Debye spheres surrounding the dust should lead to enhancement of the superdiffusion. It was also proposed in [73] that $\kappa$ is positively correlated with magnetic field fluctuations $\delta B$, which would imply negative correlation between $q$ and $\delta B$. While there is no magnetic field in PK-4, the charged dust particles are aligned to the external electric field. Thus, we conjecture that increase in electric field fluctuations should randomize the dust motion, thus, leading to classical diffusion. Finally, the negative correlation between $\kappa$ and the magnitude temperature fluctuations $\delta T^{2}$ proposed in [73] suggests a positive correlation between $q$ and $\delta T^{2}$. The role of temperature and temperature flucttuations in PK-4 will be further discussed int he next section. According to [76], the magnetic field fluctuations, $\delta B$, nearer the Sun around 6.96.96.96.99.79.79.79.7 AU, or even 43.643.643.643.647.247.247.247.2 AU, remain in a non-Gaussian, meta-equilibrium state. However, farther out between 8080808087878787 AU, where the system has had more time ($\approx 1$ year) to relax, the magnetic field fluctuations reach a Gaussian distribution suggesting an equilibrium in magnetic field fluctuations. In other words, at these distances the solar-wind relaxes back to an equilibrium after going through a transition in the diffusion regime from $q>5/3$ to $q<5/3$ between $\sim 47$ AU and $\sim 80$ AU. Similar nonequilibrium-to-equilibrium regime transitions can be explored more easily and cost-effectively in laboratory settings with a dusty plasma experiment.
@@ 6.3 VDFs and Temperature
As shown in Fig. 15, temperature decreases with increasing pressure for $T_{\|}$ and $T_{\perp}1$, aligning with the conclusions presented in [61], where it was seen that dust ’cooled off’ due to the increased role in dust neutral collision, which was calculated by use of pair-correlation functions. The temperatures $T_{\perp}2$, however, seem to show the opposite trend of increasing with increased pressure. The Tsallis coefficients $q_{v\parallel}$, never exhibit values greater than $5/3$, which suggests that there is no Lévy processes for this direction. We also notice that $q_{v\parallel}$ tends to decrease with increasing pressure (thus, improving equilibrium), while increasing current slightly increases $q_{v\parallel}$ (thus, driving the system away from equilibrium). This is reasonable as an increase in pressure causes an increase in neutral collisions, meaning that the random collision process brings the dust motion to equilibrium. Meanwhile, the current provides an electric field which affects the ion streaming and focusing surrounding the dust, which can cause local attractive interactions among dust particles and macroscopic regions of positive space charge.
@@ 7 Conclusions
The presented analysis of of PK-4 dusty plasma experiments highlights the features of anomalous dust diffusion resulting from an interplay between temperature fluctuations and anisotropies in the dust-dust interaction potential caused by the application of external electric field to this strongly coupled complex system. Notably, we used non-extensive Tsallis statistics to quantify the nonequilibrium state of the system in the directions along and across the electric field using q-Gaussian distributions. The nonequilibrium parameters $q_{p}$ and $q_{v}$ (extracted from position and velocity distributions) have distinctly different values for the directions parallel and perpendicular to the electric field, as shown in Table 2. Particle motion parallel to the electric field is superdiffusive and for several pressure-current cases, superdiffusion crosses over to a Lévy process. In the cross-field direction, the diffusion is again anomalous, but the bi-Gaussian fit best describing the histograms suggests a superposition of two processes. One of these processes is similar to a classical diffusion, while the other one is superdiffusive, but not Lévy. We attribute the directional dependence of diffusion to anisotropies in the dust-dust interaction potential and the dust temperature throughout the cloud. The former is caused by the anisotropic shape of the ion wakefield clouds surrounding the dust grains, which leads to the alignment of the dust particles into filaments. Inside a filament, macroscopic regions of positive space charge form along the direction of the electric field due to ion focusing causing local dust accelerations towards these attractors. Similarly, the cross-field motion of the dust is restrained due to a restoring force from the streaming ions within the wakefield structure. However, if a dust particle escapes the filament, it will experience large jumps across the cloud until it finds force balance within another filament (which is visible in the video data).
@@
In addition, the kinetic temperature is found to vary throughout the dust cloud as shown in Fig. 17. Analysis of subdomains within the dust clouds reveals that a single q-Gaussian fit works well even in the cross-field direction for small domain sizes. Since the $q_{v}$ extracted from velocity histograms quantifies the nonequilibrium state of the system, the domain analysis allows for a nuanced understanding of cloud properties, including the structural anisotropy, diffusion, and equilibrium state. Specifically, we conclude that the different domains within the same dust cloud are characterized by different diffusion regimes driving the global state away from equilibrium. Temperature gradients were identified for the dust in the PK-4. While all the temperatures were higher than room temperature, we propose that the observed high kinetic temperature is due to electrostatic fluctuations. Finally, we conclude that dusty plasma systems, such as those in PK-4, are excellent for studying nonequilibrium systems, anomalous diffusion and the physical origins of the two phenomena. As the mathematical description of these processes is universal, we expect that the dusty plasma experiments can be a practical means to investigate other nonequilibrium complex systems such as the solar wind.

### v3

@@ 6.2 Histogram Shape in the Cross-field direction
Figure 19: Components of the Bi-q-Gaussian illustrating two distinct populations: a) red $q_{1}=1.3$ and variance $\sigma_{1}=330$, while b) blue $q_{2}=1.63$ and $\sigma_{2}=63$. c) is the combined sum of the distributions in a) and b) while d) shows the same distributions overlapped but not added. These are the same as the $q_{v}$ values found for the cross-field velocity distribution in the 70 Pa 0.7 mA case.
The superposition of the two q-Gaussian distributions shown in Figure 19 d) shows that one diffusion process (in red) is characterized by a small peak but large variance, while the other one (in blue) has both a large peak and large tails. The variance (and the corresponding temperature) of the blue distribution, however, is much smaller. Thus, the observed diffusion in the cross-field direction can be viewed as a superposition of two processes - one with higher temperature close to equilibrium and a second one with a lower temperature but farther from equilibrium.
@@
In Space Physics, it is common to describe the non-Maxwellian velocity distributions using the kappa-distribution which is related to the q-Gaussian by $\kappa=\frac{1}{q-1}$. In [71], three physical mechanisms for generating a Kappa distribution were proposed: (i) Debye shielding, (ii) magnetic field biding, and (iii) temperature fluctuations. It was proposed that $\kappa$ has a negative correlation with the Debye number $N_{D}$ (number of particles in the Debye sphere). As $\kappa$ is inversely related to $q$, we expect a positive correlation between $N_{D}$ and $q$. In other words, increasing number of ions in the Debye spheres surrounding the dust should lead to enhancement of the superdiffusion. It was also proposed in [71] that $\kappa$ is positively correlated with magnetic field fluctuations $\delta B$, which would imply negative correlation between $q$ and $\delta B$. While there is no magnetic field in PK-4, the charged dust particles are aligned to the external electric field. Thus, we conjecture that increase in electric field fluctuations should randomize the dust motion, thus, leading to classical diffusion. Finally, the negative correlation between $\kappa$ and the magnitude temperature fluctuations $\delta T^{2}$ proposed in [71] suggests a positive correlation between $q$ and $\delta T^{2}$. The role of temperature and temperature fluctuations in PK-4 will be further discussed int he next section. According to [74], the magnetic field fluctuations, $\delta B$, nearer the Sun around 6.96.96.96.99.79.79.79.7 AU, or even 43.643.643.643.647.247.247.247.2 AU, remain in a non-Gaussian, meta-equilibrium state. However, farther out between 8080808087878787 AU, where the system has had more time ($\approx 1$ year) to relax, the magnetic field fluctuations reach a Gaussian distribution suggesting an equilibrium in magnetic field fluctuations. In other words, at these distances the solar-wind relaxes back to an equilibrium after going through a transition in the diffusion regime from $q>5/3$ to $q<5/3$ between $\sim 47$ AU and $\sim 80$ AU. Similar nonequilibrium-to-equilibrium regime transitions can be explored more easily and cost-effectively in laboratory settings with a dusty plasma experiment.
@@ 6.3 VDFs and Temperature
As shown in Figure 15, temperature decreases with increasing pressure for $T_{\|}$ and $T_{\perp}1$, aligning with the conclusions presented in [59], where it was seen that dust ’cooled off’ due to the increased role in dust neutral collision, which was calculated by use of pair-correlation functions. The temperatures $T_{\perp}2$, however, seem to show the opposite trend of increasing with increased pressure. The Tsallis coefficients $q_{v\parallel}$, never exhibit values greater than $5/3$, which suggests that there is no Lévy processes for this direction. We also notice that $q_{v\parallel}$ tends to decrease with increasing pressure (thus, improving equilibrium), while increasing current slightly increases $q_{v\parallel}$ (thus, driving the system away from equilibrium). This is reasonable as an increase in pressure causes an increase in neutral collisions, meaning that the random collision process brings the dust motion to equilibrium. Meanwhile, the current provides an electric field which affects the ion streaming and focusing surrounding the dust, which can cause local attractive interactions among dust particles and macroscopic regions of positive space charge.
@@
6.4 Energy Dissipation
The energy dissipation and thermal state of PK-4 dusty plasma clouds was studied in detail by McCabe, et al. [79]. It was found that the onset of polarity switching of the electric field in PK-4 (used for trapping the cloud in the camera’s field of view) causes an initial expansion in the dust cloud leading to an increase in thermal energy. Simulation results from that paper suggest that the onset of polarity switching causes a modification in the effective dust screening length due to reconfiguration of the dust during the expansion. The present work uses the same set of PK-4 experiments discussed in McCabe, et al., which allows for a meaningful comparison of results across the two studies. An important distinction is that McCabe, et al. used data obtained at and immediately after the onset of polarity switching, while the analysis here was performed on data collected later in the experiment. However, one of the main findings in McCabe, et al. is that an extended time (much greater than the Epstein drag decay) is required to dissipate thermal energy in these dusty plasma clouds. Thus, it is reasonable to expect that the clouds do not have enough time to come to an equilibrium state for the duration of these experiments.
This is confirmed by the statistical analysis presented here. Specifically, the velocity distributions for all cases were found to be non-Maxwellian with high energy tails, which is evident from the $q_{v}>1$ shown in Fig. 13. Here we further build on the study by McCabe et al. by identifying differences in the non-Maxwellian distributions obtained for velocities parallel versus perpendicular to the direction of the electric field. Parallel to the direction of the field, the $q_{v\parallel}$ values are only slightly greater than $1$, which suggests that the thermalization, or energy dissipation, along this direction should be close to classical. However, in the perpendicular direction, the $q_{v\perp}$ values suggest a crossover from a superdiffusive to a Lévy process. In other words, a small sub-population of the particles has an order of magnitude higher energy than the bulk particles, which leads to inefficient energy dissipation. Physically, the polarity switching of the electric field causes a modification of the ion wakefield surrounding the dust grains, which in turn leads to the observed self-organization of the dust particles into field-aligned chains. If a dust particle escapes a chain, it is likely to experience big jumps in the cross-field direction until it finds force balance as part of another chain. These jumps are clearly observable in the video data.
In addition to the onset of polarity switching, changes in the discharge conditions are another mechanism that can cause anisotropic interaction potential and slow energy dissipation in the PK-4 dusty plasma clouds. Hartmann, et al. [80] used a 2D particle-in-cell with Monte Carlo collisions discharge simulation to investigate inhomogeneities in the PK-4 plasma. It was found that, on $\mu s$-scale, the structure of the positive column is characterized by the propagation of ionization waves. This prediction was experimentally validated in the PK-4 replica at Baylor University, though for pressure and current conditions slightly higher than those considered in this study. In the presence of ionization waves, the electric field and ion/electron densities can reach amplitudes up to 10 times larger than their average values with important implications to the ion-wakefield mediated dust-dust interaction potential. Numerical work by Matthews, et al. [81] and Vermillion, et al. [82] showed that the propagation of ionization waves through the dust can modify the streaming velocity of ions and introduce time-dependent anisotropies in the ion wakefield structure surrounding each grain.
The statistical analysis conducted in this work provides quantitative experimental evidence for the presence of these anisotropies in the PK-4 ISS dusty plasmas. For all pressure-current conditions, the histogram shapes, nonextensive parameters $q_{p}$ and $q_{v}$, diffusion coefficients, and temperatures show differences when calculated for the direction along the electric field versus the cross-field direction. In addition, this anisotropy on the diffusion seems to become more pronounced at higher pressures. This can be inferred from single distribution fits of the diffusion in Figure 10 a) and b) in the $0.35mA$ and the $1mA$ case, where $D_{\parallel}$ increases with pressure increase, while $D_{\perp}$ decreases with pressure increase. Physically, the general decrease in most diffusion values in the transverse direction at higher pressure coincides with the observation that the dust alignment in filaments longitudinally improves at higher pressure, which has been confirmed by calculating pair correlation functions [83]. As the filamentary structures form, the interparticle separation and coupling strength change with direction, which in turn changes the flow of thermal energy and energy dissipation in these clouds. However, these trends are not exactly followed by the $0.7mA$ case, most notably, the $46.1Pa$ data point. While there may be particle tracking errors in the absolute particle positions data, we believe that the statistical treatment using position displacements and fits to histograms minimizes this source of errors. In addition, a measure of goodness of fit using $R^{2}$ and $\chi^{2}$ normalized by degrees of freedom, for all cases, including the $0.7mA-46.1Pa$, is typically around $R^{2}\approx 0.99$ for both parallel (single q-Gaussian Figure 10 a) and perpendicular (single q-Gaussian for Figure 10 b and bi-q-Gaussian for Fig. 10 c and d) distribution fits. The parallel direction distribution fits have normalized $\chi^{2}<10^{1}$ while normalized $\chi^{2}\approx 10^{2}$ for single q-Gaussian perpendicular fits and $\chi^{2}\approx 10^{1}$ in the bi-q-Gaussian fits, which is expected due to better fitting of the tail behavior. Thus instead of fitting error, the probable source of the disagreement in the $46.1Pa$ $0.7mA$ data point is the difference in dust densities for different plasma conditions. Re-examination of the data from the PK-4 plasma glow camera, we could see that for all currents, the $46.1Pa$ experiments had one large cloud at the center of the discharge tube, while for the $28.5Pa$ and the $70.5Pa$ cases, there were multiple smaller dust clouds forming around plasma striations throughout the discharge tube. Since the number of particle dispenser shakes was kept constant, it is reasonable to expect that the difference in size and number of clouds will lead to a difference in dust density, which in turn, affects the calculated statistical properties. In a follow-up experiment (not discussed here), the plasma conditions were kept constant, while the dust density of the same cloud was varied through cloud compressions. These results will be a subject of a follow-up paper.
It has been demonstrated numerically [81] that the formation of filamentary structures in PK-4 is enhanced by the presence of ionization waves. Since the PK-4 experiment on the ISS is not equipped with a high-speed video camera, we do not have direct experimental evidence that these ionization waves occur for all examined pressure-current conditions. However, the statistical analysis presented here provides a way to quantify the possible effect of ionization waves through the observed anisotropies in the dust diffusion. The present findings will benefit from comparison to further discharge simulations and experimental studies using the on-ground PK-4 replicas to better understand the parameter space where the ionization waves should be expected to dominate.
@@ 7 Conclusions
The presented analysis of of PK-4 dusty plasma experiments highlights the features of anomalous dust diffusion resulting from an interplay between temperature fluctuations and anisotropies in the dust-dust interaction potential caused by the application of external electric field to this strongly coupled complex system. Notably, we used non-extensive Tsallis statistics to quantify the nonequilibrium state of the system in the directions along and across the electric field using q-Gaussian distributions. The nonequilibrium parameters $q_{p}$ and $q_{v}$ (extracted from position and velocity distributions) have distinctly different values for the directions parallel and perpendicular to the electric field, as shown in Table 2. Particle motion parallel to the electric field is superdiffusive and for several pressure-current cases, superdiffusion crosses over to a Lévy process. In the cross-field direction, the diffusion is again anomalous, but the bi-Gaussian fit best describing the histograms suggests a superposition of two processes. One of these processes is similar to classical diffusion, while the other one is superdiffusive, but not Lévy. We attribute the directional dependence of diffusion to anisotropies in the dust-dust interaction potential and the dust temperature throughout the cloud. The former is caused by the anisotropic shape of the ion wakefield clouds surrounding the dust grains, which leads to the alignment of the dust particles into filaments. Inside a filament, macroscopic regions of positive space charge form along the direction of the electric field due to ion focusing causing local dust accelerations towards these attractors. Similarly, the cross-field motion of the dust is restrained due to a restoring force from the streaming ions within the wakefield structure. However, if a dust particle escapes the filament, it will experience large jumps across the cloud until it finds force balance within another filament (which is visible in the video data). We hypothesize that the higher parallel diffusion causes more energy exchange between particles, thus bringing the system to equilibrium. This is seen by $q_{v\parallel}$ being only slightly greater than $1$ while the parallel diffusion is large. Also, $q_{v\perp}$ is relatively greater than $1$ and the perpendicular diffusion is much smaller. This is also supported by the trends that exist between the parallel diffusion and $q_{v\parallel}$ with increasing pressure, which are inversely proportional. The same is true of perpendicular diffusion and $q_{v\perp}$ also being inversely proportional, which may suggest that increased diffusion contributes to increasing the equilibrium.
@@
In addition, the kinetic temperature is found to vary throughout the dust cloud as shown in Figure 17. Analysis of subdomains within the dust clouds reveals that a single q-Gaussian fit works well even in the cross-field direction for small domain sizes. Since the $q_{v}$ extracted from velocity histograms quantifies the nonequilibrium state of the system, the domain analysis allows for a nuanced understanding of cloud properties, including the structural anisotropy, diffusion, and equilibrium state. Specifically, we conclude that the different domains within the same dust cloud are characterized by different diffusion regimes driving the global state away from equilibrium. Temperature gradients were identified for the dust in the PK-4. While all the temperatures were higher than room temperature, we propose that the observed high kinetic temperature is due to electrostatic fluctuations. Finally, we conclude that dusty plasma systems, such as those in PK-4, are excellent for studying nonequilibrium systems, anomalous diffusion and the physical origins of the two phenomena. As the mathematical description of these processes is universal, we expect that the dusty plasma experiments can be a practical means to investigate other nonequilibrium complex systems such as the solar wind.